triangle yzx given has 3 congruent angles 60 degrees top leg = 8. another leg = 3x-y, last leg is x + 3y + 2. find x and y. This is section 4.4 masters level b p.71 using triangle congruence holt 2004 problem 5. I have the answer which is x = 3 y = 1 but have no idea how they get there and I have to teach my son. Thanks for any help in understanding its been 30 something years.
Answer by scott8148(6628)

the triangle has three equal angles, so it also has three equal sides

8 = 3x + y = x + 3y + 2

3x + y = 8 ___ y = 8 - 3x

x + 3y + 2 = 8 ___ x + 3y = 6

substituting ___ x + 3(8 - 3x) = 6 ___ x + 24 - 9x = 6

24 - 8x = 6 ___ 18 = 8x ___ 9/4 = x

substituting ___ y = 8 - 3(9/4) = 5/4